# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v2

@@ 11. Large Legal Corpora
Training LLMs for legal NLP requires extensive legal corpora that are transparent in sourcing, safeguard privacy, and minimise bias to ensure fairness and accuracy. These corpora serve as the foundation for developing models capable of handling diverse legal tasks. To assess model performance, evaluation benchmarks provide structured datasets and standardised metrics for tasks such as judgement prediction, QA, case retrieval, and entailment. Together, large legal corpora and evaluation benchmarks try to support the advancement of reliable legal AI applications.
@@ 11.1. Evaluation benchmarks
Zheng et al. (2021) introduce the CaseHOLD dataset, a novel benchmark for evaluating NLP models in the legal domain, designed to address the challenge of identifying the legal holdings from case texts. The dataset contains over 53,000 multiple choice questions derived from U.S. case law citations, where each question requires the identification of the correct holding from a set of potential answers. This task, simulating a fundamental lawyering skill taught in law school, involves contextual understanding and application of legal rules to factual situations. CaseHOLD is aimed at enhancing model training by focusing on semantic matching and the ability to discern legal principles. The dataset is structured to provide a challenging yet accessible resource for NLP researchers, with a clear focus on promoting deeper understanding and application of legal rules in automated systems. It utilises a format where a cited text serves as a prompt with five answer options—one correct holding and four closely related incorrect holdings—to refine the models’ abilities to reflect legal reasoning.
@@
Chalkidis et al. (2022a) introduce the Legal General Language Understanding Evaluation (LexGLUE) benchmark, a comprehensive suite of datasets aimed at assessing the capabilities of NLP models across various legal tasks. The benchmark covers datasets, such as ECtHR (Chalkidis et al., 2019a), SCOTUS [^9] https://www.supremecourt.gov, EUR-LEX, LEDGAR (Tuggener et al., 2020), UNFAIR-ToS (Lippi et al., 2019), and CaseHOLD (Zheng et al., 2021), each chosen for its complexity, relevance, and need for legal expertise. These datasets cover a range of tasks from multi-label and multi-class classification to multiple-choice questions and are split chronologically into training, development, and test sets to provide standardised evaluation metrics. For instance, ECtHR datasets focus on violations of European Convention of Human Rights provisions, SCOTUS database classifies U.S. Supreme Court opinions by legal issues, EUR-LEX database involves labelling EU laws with EuroVoc concepts, LEDGAR classifies provisions of U.S. contracts, UNFAIR-ToS identifies unfair terms in online service agreements, and CaseHOLD involves QA about legal rulings.
@@
Chalkidis et al. (2022b) present FairLex, a benchmark suite consisting of four legal datasets—ECtHR (Chalkidis et al., 2019a), SCOTUS, FSCS, and CAIL—that address the fairness of NLP applications across diverse legal jurisdictions and languages, including English, German, French, Italian, and Chinese. These datasets, curated from European Council, USA, Switzerland, and China, cover various legal tasks such as LJP, issue area classification, and crime severity prediction, aiming to test the performance and fairness of LMs in recognising and classifying legal texts. FairLex focuses on ensuring demographic, regional, and legal topic fairness by analysing attributes, such as gender, age, region of origin, and legal areas within cases. Each dataset in FairLex provides a substantial number of cases, systematically divided into training, development, and test sets, and includes detailed attributes, such as the defendant state in ECtHR, decision direction in SCOTUS, legal areas in FSCS, and demographic details in CAIL.
@@
Rabelo et al. (2022) summarise the 8th Competition on Legal Information Extraction and Entailment (COLIEE 2021), which featured five tasks across case and statute law, engaging participants from various teams to apply diverse NLP approaches. The competition tasks included case law retrieval and entailment, as well as statute law retrieval and entailment with and without prior retrieved data. Specifically, task 1 focused on extracting relevant supporting cases from a corpus, while task 2 involved identifying paragraphs from cases that entail a given new case fragment. For statute law, tasks 3 and 4 entailed retrieving and answering questions based on civil code statutes, with task 5 challenging participants to answer without pre-retrieved statutes. The datasets used varied in complexity, from 4415 case files in task 1 with a need to identify noticed cases without relying on citations, to the civil code-based tasks 3, 4, and 5 which adapted to recent legal revisions in Japanese law and excluded untranslated parts, reflecting the ongoing evolution and challenge in legal NLP applications.
@@
Barale et al. (2023) present AsyLex, a pioneering dataset tailored for refugee law applications, featuring 59,112 documents from Canadian refugee status determinations spanning from 1996 to 2022. This dataset is designed to enhance the capabilities of NLP models in legal research by providing 19,115 gold-standard human-annotated and 30,944 inferred labels for entity extraction and LJP. Key contributions include anonymizing decision documents, employing a robust annotation methodology, and creating datasets for specific NLP tasks, such as entity extraction and judgement prediction. This rich corpus, with detailed annotations across 22 categories, supports complex legal NLP tasks, thereby filling the gap in resources for the legal domain.
Niklaus et al. (2023) introduce LEXTREME, a multilingual benchmark specifically designed to evaluate LMs on legal NLP tasks, a critical step given the unique challenges of legal language. Surveying legal NLP literature from 2010 to 2022, they curate 11 datasets spanning 24 languages and cover a variety of legal domains, employing datasets that only involve human-annotated texts or those with annotations derived through clear methodological frameworks. They introduce two aggregate scores to facilitate fair comparison across models: the dataset aggregate score and the language aggregate score, revealing a performance correlation with model size on LEXTREME. The benchmark consists of three task types: single label text classification, multi label text classification, and NER, using existing splits for training, validation, and testing when available, or creating random splits otherwise. This effort marks an advancement in testing NLP capabilities across a diverse range of legal documents and languages.
@@
Park and James (2023) explore the creation of a Natural Language Inference dataset within the legal domain, focusing on criminal court verdicts in Korean. Their methodology includes the innovative use of adversarial hypothesis generation to challenge annotators and enhance the robustness of the dataset, supported by visual tools for hypothesis network construction. The data collection involves extracting context from verdicts and augmenting it using Easy Data Augmentation (Wei and Zou, 2019) techniques and round-trip translation to generate a dataset for training and testing Natural Language Inference models. The study highlights issues such as annotators’ limited domain knowledge and challenges in handling long contexts but provides solutions, such as targeted data collection and the use of gamification to boost annotator engagement and productivity.
@@
Goebel et al. (2024) summarise COLIEE 2023, featuring four tasks across case and statute law with participation from ten different teams engaging in multiple tasks. task 1 involves legal case retrieval, requiring participants to extract supporting cases from a corpus, and task 2 focuses on legal case entailment, identifying paragraphs that entail aspects of a new case. Task 3 and 4, based on Japanese civil code statutes from the bar exam, involve retrieving relevant articles and verifying statements, respectively. The competition leverages a dataset of over 5,700 case law files and introduces new query cases and test questions sourced from recent bar exams, testing the efficacy of different teams’ approaches in handling complex legal texts and hypotheses in a controlled competitive environment.
@@
Östling et al. (2024) introduce the Cambridge Law Corpus (CLC), a legal dataset featuring 258,146 cases from UK courts, dating from the 16th century to the present. The corpus includes raw text and metadata across various court types, and is structured in XML format for ease of use and annotated for case outcomes in a subset of 638 cases. Additionally, the CLC is supported by a Python library for data manipulation and ML applications.
@@ 11.2. Large corpora for pre-training
Henderson et al. (2022) introduce the ‘Pile of Law’, the first and an important large corpus in the legal domain, containing a 256 GB dataset of open-source English-language legal and administrative data. This dataset includes contracts, court opinions, legislative records, and administrative rules, curated to explore data sanitation norms across legal and administrative settings and serve as a tool for pre-training legal-domain LMs. They emphasise the legal norms governing privacy and toxicity filtering, detailing how the dataset reflects these norms through built-in filtering mechanisms in the collected data, which include court filings, legal analyses, and government publications. By analysing how legal and administrative entities handle sensitive information and potentially offensive content, the paper provides actionable insights for researchers to improve content filtering practices before pre-training LLMs, thereby enhancing the ethical use of NLP in legal applications.
Niklaus et al. (2024) present the MultiLegalPile, the largest open-source multilingual legal corpus available, totalling 689 GB and spanning 17 jurisdictions across 24 languages. This extensive dataset is designed to facilitate training of LLMs within the legal domain, featuring diverse legal text types including case law, legislation, and contracts, predominantly in English due to the integration of the ‘Pile of Law’ (Henderson et al., 2022) dataset. Through careful regex-based filtering from the mC4 corpus and manual reviews, the team ensures high precision in legal content selection.
@@ 12. Legal Language Models and Methods for Legal Domain Adaptation
In the fast-moving field of NLP, LLMs have become a key tool for processing and understanding large amounts of unstructured text data. These models, initially trained on broad datasets, such as Wikipedia, have shown great skill across various language tasks. Building on this success, the legal technology community is increasingly interested in using these powerful models for legal NLP applications. This involves adapting these general-domain models to legal texts and further training them on specialised legal documents. Such efforts aim to reduce the domain gap and customise the models to better understand the complex language used in legal documents. In this section, we will explore how these models are being adapted and applied within the legal domain to enhance legal NLP applications.
In this section, following the methodology of this survey, we studied all peer-reviewed LMs or methods. However, due to the challenges present in the legal domain, there are many legal LMs that have not undergone peer review. Given the scarcity of adequate peer-reviewed resources, our research has focused on the investigation of, in order of priority, the peer-reviewed sources, then the most well-known and widely used non-peer-reviewed legal LMs. Despite their lack of formal peer review, these models have gained considerable attention and usage in the field.
@@ 12.1. Language Models
Chalkidis et al. (2020) present an in-depth analysis of applying BERT, a pre-trained LM, in the legal domain, showcasing the need for domain-specific adaptation to enhance performance on legal NLP tasks. They explore three strategies: using standard BERT directly, further pre-training on legal corpora, and pre-training from scratch with legal-specific data. Their study found that both further pre-training and pre-training from scratch generally outperform the use of BERT directly. They introduce legal-bert, a specialised family of models optimised for legal text, which includes versions for varied computational capacities and demonstrates competitive performance with a lower environmental impact.
@@
In the development of specialised NLP tools for Arabic legal texts, a model specifically tailored to the unique linguistic features of Arabic jurisprudence was designed, introducing AraLegal-BERT (Al-qurishi et al., 2022) midway through this innovation. This model enhances NLP applications within the legal field by adapting BERT technology to Arabic’s specific content needs, involving pre-training BERT from scratch using a broad range of legal documents, including legislative materials and contracts.
@@
Shi et al. (2024) develop Legal-LM, a specialised LM tailored for Chinese legal consulting, enhanced with a KG to address domain-specific challenges such as data veracity and non-expert user interaction. The framework involves several steps: extensive pre-training on a rich corpus of legal texts integrated with a legal KG, keyword extraction and Direct Preference Optimisation to refine responses, and the use of an external legal knowledge base for data retrieval and response validation. This multi-faceted approach ensures that Legal-LM not only comprehends complex legal language but also generates precise and user-aligned legal advice.
@@ 12.2. Methods for Improving In-Domain Adaptability of Legal Language Models
Mamakas et al. (2022) explore strategies for adapting pre-trained transformers to cope with the challenges of long legal texts within the LexGLUE benchmark, focusing on extending input capabilities and enhancing efficiency. They modify Longformer (Beltagy et al., 2020), originally extending up to 4,096 sub-words, to process up to 8,192 sub-words by reducing local attention window size and incorporating a global token at the end of each paragraph to facilitate information flow across longer texts. Additionally, they adapt legal-bert to employ TF-IDF representations to manage longer documents, introducing variants, such as TF-IDF-SRT-LegalBERT, which deduplicates and sorts sub-words by TF-IDF scores, and TF-IDF-EMB-LegalBERT, which incorporates a TF-IDF embedding layer. These adaptations aim to combine the robust capabilities of transformers with the practical requirements of handling extensive legal documents, surpassing the performance of traditional linear classifiers while maintaining computational efficiency.
@@ 13. Open Research Challenges
Despite researchers’ efforts in the this interdisciplinary field and extensive advancements in AI techniques, Open Research Challenges (ORCs) still exist. In this section, we identify the ORCs and provide advice to overcome these challenges.
@@
Bias and fairness are crucial concerns in the field of AI, especially at the intersection with the legal domain where decisions can deeply impact individuals’ lives. The scarcity of unbiased data in legal domains such as case law complicates the training of AI models, as these models often learn from historical decisions that may reflect existing human biases (Teichman et al., 2023; Edmond and Martire, 2019). This reliance on biased datasets can lead to unfair and biased outcomes in classification and prediction tasks. Addressing these issues is critical to ensure that AI-driven legal decisions uphold the standards of impartiality and fairness required for justice.
ORC2: Privacy Concern
Privacy concerns in legal NLP are critical, as models often handle highly sensitive documents such as court records (Yin and Habernal, 2022). Beyond basic anonymization, advanced techniques, such as differential privacy add statistical noise during training to ensure that individual data points have minimal influence on the model’s output. Adversarial training further bolsters privacy by simulating potential attacks and guiding the model to suppress identifiable attributes in its representations. Privacy-preserving fine-tuning methods—such as applying differential privacy during fine-tuning, federated learning, or knowledge distillation—allow models to adapt to new legal tasks without exposing sensitive data. However, implementing these techniques requires careful calibration, as excessive privacy constraints may degrade model accuracy. This ongoing challenge underscores the need for continued research into privacy-preserving methods tailored to the unique demands of the legal domain.
@@
Interpretability and explainability are crucial across various applications in legal NLP, yet these aspects remain underexplored in many studies. The ability to trace and comprehend the decision-making process of AI systems is essential for identifying and mitigating biases. Transparent and understandable AI systems help build trust and ensure they are used responsibly, which is particularly important in legal contexts where decisions can significantly impact people’s lives. Improving these aspects of AI models is necessary to their ethical use, ensuring they meet the high standards of fairness required in legal proceedings.
@@
Annotation quality and transparency are critical challenges in legal NLP. A growing discussion in the community focuses on descriptive versus prescriptive annotation (Röttger et al., 2022). Descriptive annotation captures the full spectrum of annotator subjectivity, preserving diverse interpretations of legal texts and proving valuable for interpretative tasks such as legal question answering or contract analysis. In contrast, prescriptive annotation enforces a single, consistent standard that is essential for tasks requiring strict adherence to legal norms, such as LJP or statute classification. The choice between these paradigms depends on the specific task, domain context, and intended downstream application. Equally, transparent annotation practices are rarely documented in legal NLP studies, with many reports omitting details about annotators’ backgrounds and the procedures employed. Comprehensive documentation of the annotation process is important for assessing dataset quality, ensuring fairness, and building trust in legal NLP systems. Ultimately, explicit decisions regarding both the annotation paradigm and transparency measures are necessary to produce reliable, suitable datasets.
@@
One of the key ORCs in legal NLP is the scarcity of annotated data, which limits the development of robust models due to the high cost and expertise required for labelling complex legal texts. Data augmentation can offer a promising approach to mitigate this issue, yet it introduces its own set of challenges. The complexity of legal texts complicates the application of general augmentation techniques. Ensuring the quality of augmented data remains critical, with efforts focusing on preserving semantic and syntactic integrity to prevent models from learning incorrect patterns (Perçin et al., 2022). Additionally, the effectiveness of augmentation varies across models, necessitating tailored strategies, while scalability concerns arise with computationally intensive methods (Ghosh et al., 2023). The predominant focus on English limits applicability to multilingual legal contexts, and in extremely low-resource scenarios, there’s a risk of overfitting to augmented data (Ghosh et al., 2023). Future research should prioritise developing domain-specific augmentation techniques that maintain legal text nuances, explore flexible approaches, enhance scalability, and extend methods to support diverse languages, thereby addressing these persistent challenges in legal NLP.
@@
In legal NLP, enhancing multilingual capabilities remains an underdeveloped area. While efforts, such as MultiLegalPile (Niklaus et al., 2024) have begun to address this, there remains a gap in research for many languages, including but not limited to Persian and Arabic. These limitations restricts the application of legal NLP across diverse legal systems worldwide, which is important for broader accessibility. Multilingual capabilities introduce unique challenges for legal NLP models, primarily due to the distinct linguistic structures of each language, which often require extensive fine-tuning to ensure accuracy and relevancy in legal contexts. Furthermore, each legal system possesses its own set of terms and document standards, which can vary dramatically from one language to another. Therefore, expanding research into these and other underserved languages is essential for making NLP tools universally applicable.
ORC7: Ontology
The use of ontologies in the legal domain is relatively sparse, yet it holds considerable potential to enhance the robustness of AI methodologies. Ontology or KG can also enable the AI models to draw accurate inferences regarding the relationship between the terms and thereby better understand and process complicated legal texts. This approach could advance the capability of AI systems to handle complex legal reasoning and decision-making processes. However, utilising ontology in legal NLP faces unique challenges. The complexity of legal language and the concept of ‘open texture’, where the meaning of legal terms can evolve over time, complicate the creation of static ontological models (Mommers, 2010). Legal ontologies must be dynamic, reflecting changes in law and its interpretation over time. Additionally, the integration of real-world and legal concepts within ontologies presents further complexity, as it requires accommodating both legal terms and their relevant real-world contexts (Mommers, 2010).

### v3

@@ 11. Large Legal Datasets
Training LLMs for legal NLP requires extensive legal corpora that are transparent in sourcing, safeguard privacy and minimise bias to ensure fairness and accuracy. These corpora serve as the foundation for developing models capable of handling diverse legal tasks. To assess model performance, evaluation benchmarks provide structured datasets and standardised metrics for tasks such as judgement prediction, QA, case retrieval and entailment. Together, large legal corpora and evaluation benchmarks can support the advancement of reliable legal AI applications.
Table 4. Overview of datasets and their usage across different legal NLP tasks and cited studies.
Task Datasets Identified Studies
LJP CAIL datasets by Xiao et al. (2018) Yang et al. (2019); Zhong et al. (2018); Yang et al. (2019); Zhong et al. (2020a); Xu et al. (2020); Feng et al. (2022); Tong et al. (2024); Zhang et al. (2023a)
LTC LEDGAR by Tuggener et al. (2020) Mamooler et al. (2022)
EURLEX57K by Chalkidis et al. (2019b) Song et al. (2022)
NER LegalNERo by Păis et al. (2021) Smădu et al. (2022)
German LER by Leitner et al. (2020)
LAM ECHR by Poudyal et al. (2020) Zhang et al. (2023b)
Demosthenes by Santin et al. (2023) Zhang et al. (2023b); Santin et al. (2023)
@@ 11.1. Evaluation benchmarks
Zheng et al. (2021) introduce CaseHOLD, a novel benchmark for evaluating NLP models in the legal domain, designed to address the challenge of identifying the legal holdings from case texts. The dataset contains over 53,000 multiple-choice questions derived from the US case law citations, where each question requires the identification of the correct holding from a set of potential answers. This task, simulating a fundamental skill taught in law school, involves contextual understanding and application of legal rules to factual situations. CaseHOLD is aimed at enhancing model training by focusing on semantic matching and the ability to discern legal principles. The dataset is structured to provide a challenging yet accessible resource for NLP researchers, with a clear focus on promoting deeper understanding and application of legal rules in automated systems. Each item presents a cited passage as a prompt followed by one correct holding and four closely related incorrect options, encouraging models to demonstrate genuine legal reasoning.
@@
Chalkidis et al. (2022a) introduce the Legal General Language Understanding Evaluation (LexGLUE) benchmark, a comprehensive suite of datasets aimed at assessing the capabilities of NLP models across various legal tasks. The benchmark covers datasets, such as ECHR (Chalkidis et al., 2019a), SCOTUS, [^8] https://www.supremecourt.gov EUR-Lex, LEDGAR (Tuggener et al., 2020), UNFAIR-ToS (Lippi et al., 2019) and CaseHOLD (Zheng et al., 2021), each chosen for its complexity, relevance, and need for legal expertise. These datasets cover a range of tasks from multi-label and multi-class classification to multiple-choice questions and are split chronologically into training, development and test sets to provide standardised evaluation metrics. For instance, ECHR datasets focus on violations of European Convention of Human Rights provisions; theSCOTUS database classifies the US Supreme Court opinions by legal issues; the EUR-Lex database involves labelling EU laws with EuroVoc concepts; LEDGAR classifies provisions of the US contracts; UNFAIR-ToS identifies unfair terms in online service agreements; and CaseHOLD involves QA about legal rulings.
@@
Chalkidis et al. (2022b) introduce FairLex, a multilingual benchmark suite of four legal datasets: ECHR (Chalkidis et al., 2019a), SCOTUS, FSCS and CAIL. The suite evaluates fairness in NLP across several jurisdictions (Europe, the United States, Switzerland, and China) and five languages (English, German, French, Italian, and Chinese). Each dataset is aligned with a specific legal task: ECHR violation prediction for ECHR; issue-area classification for SCOTUS; case-approval prediction for FSCS; and crime-severity prediction for CAIL. All datasets are chronologically split into training, development and test sets. FairLex supports demographic, regional and topical fairness analysis by recording sensitive attributes, including defendant state in ECHR, decision direction in SCOTUS, legal area in FSCS, and gender and region of origin in CAIL.
@@
Rabelo et al. (2022) summarise the 8th Competition on Legal Information Extraction and Entailment (COLIEE 2021), which featured five tasks across case and statute law, engaging participants from various teams to apply diverse NLP approaches. The competition tasks included case law retrieval and entailment, as well as statute law retrieval and entailment with and without prior retrieved data. Specifically, task 1 focused on extracting relevant supporting cases from a corpus, while task 2 involved identifying paragraphs from cases that entail a given new case fragment. For statute law, tasks 3 and 4 entailed retrieving and answering questions based on civil code statutes, with task 5 challenging participants to answer without pre-retrieved statutes. The datasets used varied in complexity, from 4,415 case files in task 1 with a need to identify noticed cases without relying on citations, to the civil code-based tasks 3, 4 and 5 which adapted to recent legal revisions in Japanese law and excluded untranslated parts, reflecting the ongoing evolution and challenge in legal NLP applications.
@@
Barale et al. (2023) present AsyLex, a pioneering dataset tailored for refugee law applications, featuring 59,112 documents from Canadian refugee status determinations spanning from 1996 to 2022. This dataset is designed to enhance the capabilities of NLP models in legal research by providing 19,115 gold-standard human-annotated and 30,944 inferred labels for entity extraction and LJP. Key contributions include anonymizing decision documents, employing a robust annotation methodology and creating datasets for specific NLP tasks, such as entity extraction and judgement prediction.
Niklaus et al. (2023) present LEXTREME, a multilingual benchmark for evaluating LMs on legal NLP tasks. Drawing on legal NLP research published between 2010 and 2022, they curate 11 human-annotated datasets spanning 24 languages and multiple legal domains. To ensure fair comparison across models, the authors introduce two aggregate metrics—the dataset aggregate score and the language aggregate score—and show that performance on LEXTREME rises with model size. The benchmark covers three task types: single-label text classification, multi-label text classification and NER. Where possible, it preserves existing train, development and test splits, otherwise creating random splits.
@@
Park and James (2023) explore the creation of a Natural Language Inference (NLI) dataset within the legal domain, focusing on criminal court verdicts in Korean. Their methodology includes the innovative use of adversarial hypothesis generation to challenge annotators and enhance the robustness of the dataset, supported by visual tools for hypothesis network construction. The data collection involves extracting context from verdicts and augmenting it using Easy Data Augmentation (Wei and Zou, 2019) techniques and round-trip translation to generate a dataset for training and testing NLI models. The study highlights issues such as annotators’ limited domain knowledge and challenges in handling long contexts but provides solutions, such as targeted data collection and the use of gamification to boost annotator engagement and productivity.
@@
Goebel et al. (2024) summarise COLIEE 2023, featuring four tasks across case and statute law with participation from ten different teams engaging in multiple tasks. task 1 involves legal case retrieval, requiring participants to extract supporting cases from a corpus and task 2 focuses on legal case entailment, identifying paragraphs that entail aspects of a new case. Task 3 and 4, based on Japanese civil code statutes from the bar exam, involve retrieving relevant articles and verifying statements, respectively. The competition leverages a dataset of over 5,700 case law files and introduces new query cases and test questions sourced from recent bar exams, testing the efficacy of different teams’ approaches in handling complex legal texts and hypotheses in a controlled competitive environment.
@@
Östling et al. (2024) introduce the Cambridge Law Corpus (CLC), a legal dataset featuring 258,146 cases from UK courts, dating from the 16th century to the present. The corpus includes raw text and metadata across various court types and is structured in XML format for ease of use and annotated for case outcomes in a subset of 638 cases. Additionally, the CLC is supported by a Python library for data manipulation and ML applications.
@@ 11.2. Large corpora for pre-training
Henderson et al. (2022) introduce the “Pile of Law,” the first at-scale legal text collection, containing a 256 GB dataset of open-source English-language legal and administrative data. This dataset includes contracts, court opinions, legislative records, and administrative rules, curated to explore data sanitation norms across legal and administrative settings and serve as a tool for pre-training legal LLMs. They emphasise the legal norms governing privacy and toxicity filtering, detailing how the dataset reflects these norms through built-in filtering mechanisms in the collected data, which include court filings, legal analyses and government publications. By analysing how legal and administrative entities handle sensitive information and potentially offensive content, the paper provides actionable insights for researchers to improve content filtering practices before pre-training LLMs, thereby enhancing the ethical use of NLP in legal applications.
Niklaus et al. (2024) present the MultiLegalPile, the largest open-source multilingual legal corpus available, totalling 689 GB and spanning 17 jurisdictions across 24 languages. This extensive dataset is designed to facilitate training of LLMs within the legal domain, featuring diverse legal text types including case law, legislation and contracts, predominantly in English due to the integration of the “Pile of Law” (Henderson et al., 2022) dataset. Through careful regex-based filtering from the mC4 corpus and manual reviews, the team ensures high precision in legal content selection.
@@ 12. Legal Language Models and Methods for Legal Domain Adaptation
In the fast-moving field of NLP, LLMs have become a key tool for processing and understanding large amounts of unstructured text data. These models, initially trained on broad datasets such as Wikipedia, have shown great skill across various language tasks. Building on this success, the legal technology community is increasingly interested in using these powerful models for legal NLP applications. This involves adapting these general-domain models to legal texts and further training them on specialised legal documents. Such efforts aim to reduce the domain gap and customise the models to better understand the complex language used in legal documents. In this section, we will explore how these so-called “foundation” models are being adapted and applied within the legal domain to enhance legal NLP applications.
Following the methodology of this survey, this section studies all peer-reviewed LMs or related methods. However, due to the challenges present in the legal domain, there are many legal LMs that have not undergone peer review. Given the scarcity of adequate peer-reviewed resources, our research has focused on the investigation of, in order of priority, the peer-reviewed sources, then the most well-known and widely used non-peer-reviewed legal LMs. Despite their lack of formal peer review, these models have gained considerable attention and usage in the field, and some are expected to be published in peer-reviewed venues in the future.
@@ 12.1. Language Models
Chalkidis et al. (2020) present an in-depth analysis of applying BERT in the legal domain, showcasing the need for domain-specific adaptation to enhance performance on legal NLP tasks. They explore three strategies: using standard BERT directly, further pre-training a BERT model on legal corpora, and pre-training from scratch with legal-specific data. Their study found that both further pre-training and pre-training from scratch generally outperform the use of BERT directly. They introduce legal-bert, which includes versions for varied computational capacities and demonstrates competitive performance with a lower environmental impact.
@@
In the development of specialised NLP tools for Arabic legal texts, a model specifically tailored to the unique linguistic features of Arabic jurisprudence was designed called AraLegal-BERT (Al-qurishi et al., 2022). This model enhances NLP applications within the legal field by adapting BERT technology to Arabic’s specific content needs, involving pre-training BERT from scratch using a broad range of legal documents, including legislative materials and contracts.
@@
Shi et al. (2024) develop Legal-LM, a specialised LM tailored for Chinese legal consulting, enhanced with a KG to address domain-specific challenges such as data veracity and non-expert user interaction. The framework involves several steps: extensive pre-training on a rich corpus of legal texts integrated with a legal KG, keyword extraction and Direct Preference Optimisation to refine responses and the use of an external legal knowledge base for data retrieval and response validation. This multi-faceted approach ensures that Legal-LM not only comprehends complex legal language but also generates precise and user-aligned legal advice.
@@ 12.2. Methods for Improving In-Domain Adaptability of Legal Language Models
Mamakas et al. (2022) explore strategies for adapting pre-trained transformers to cope with the challenges of long legal texts within the LexGLUE benchmark, focusing on extending input capabilities and enhancing efficiency. They modify Longformer (Beltagy et al., 2020), originally extending up to 4,096 subwords, to process up to 8,192 subwords by reducing local attention window size and incorporating a global token at the end of each paragraph to facilitate information flow across longer texts. Additionally, they adapt legal-bert to employ TF-IDF representations to manage longer documents, introducing variants, such as TF-IDF-SRT-LegalBERT, which deduplicates and sorts subwords by TF-IDF scores; and TF-IDF-EMB-LegalBERT, which incorporates a TF-IDF embedding layer. These adaptations aim to combine the robust capabilities of transformers with the practical requirements of handling extensive legal documents, surpassing the performance of traditional linear classifiers while maintaining computational efficiency.
@@ 13. Open Research Challenges
Despite researchers’ efforts in the this interdisciplinary field and extensive advancements in AI techniques, a number of open research challenges (ORCs) still exist. In this section, we identify the key ORCs, and provide advice and directions for future work to overcome these challenges.
@@
Bias and fairness are crucial concerns in the field of AI, especially at the intersection with the legal domain where decisions can deeply impact individuals’ lives. The scarcity of unbiased data in legal domains such as case law complicates the training of AI models, as these models often learn from historical decisions that may reflect existing human biases (Teichman et al., 2023; Edmond and Martire, 2019). This reliance on biased datasets can, in turn, lead to unfair and biased outcomes in downstream classification and prediction tasks. Addressing these issues is critical to ensure that AI-driven legal decisions uphold the standards of impartiality and fairness required for justice.
ORC2: Privacy Concerns
Privacy concerns in legal NLP are critical, as models often handle highly sensitive documents such as court records (Yin and Habernal, 2022). Beyond basic anonymization, advanced techniques – such as differential privacy – add statistical noise during training to ensure that individual data points have minimal influence on the model’s output. Adversarial training further bolsters privacy by simulating potential attacks and guiding the model to suppress identifiable attributes in its representations. Privacy-preserving fine-tuning methods – such as applying differential privacy during fine-tuning, federated learning or knowledge distillation – allow models to adapt to new legal tasks without exposing sensitive data. However, implementing these techniques requires careful calibration, as excessive privacy constraints may degrade model accuracy. This challenge motivates the need for continued research into privacy-preserving NLP methods tailored to the unique demands of the legal domain.
@@
The ability to interpret and explain the outputs of AI-powered systems is of critical importance across various applications in legal NLP, yet these aspects remain underexplored in many contexts. Relating to ORC1, the ability to trace and comprehend the decision-making process of AI systems is essential for identifying and mitigating biases. Transparent and understandable AI systems help build trust and ensure they are used responsibly, which is particularly important in legal contexts where decisions significantly impact people’s lives. Improving these aspects of AI models is necessary to their ethical use, ensuring they meet the high standards of fairness required in legal proceedings. In turn, this will promote wider adaption, allowing the promises of these systems – increased productivity, fairness, and reduced costs – to be realised in practice.
@@
Annotation quality and transparency are critical challenges in legal NLP. A growing discussion in the community focuses on descriptive versus prescriptive annotation (Röttger et al., 2022). Descriptive annotation captures the full spectrum of annotator subjectivity, preserving diverse interpretations of legal texts and proving valuable for interpretative tasks such as LQA or contract analysis. In contrast, prescriptive annotation enforces a single, consistent standard that is essential for tasks requiring strict adherence to legal norms, such as LJP or statute classification. The choice between these paradigms depends on the specific task, domain context and intended downstream application. Equally, transparent annotation practices are rarely documented in legal NLP studies, with many reports omitting details about annotators’ backgrounds and the procedures employed. Comprehensive documentation of the annotation process is important for assessing dataset quality, ensuring fairness and building trust in legal NLP systems. Ultimately, explicit decisions regarding both the annotation paradigm and transparency measures are necessary to produce reliable, suitable datasets.
@@
One of the key ORCs in legal NLP is the scarcity of annotated data, which limits the development of robust models due to the high cost and expertise required for labelling complex legal texts. Data augmentation can offer a promising approach to mitigate this issue, yet it introduces its own set of challenges. The complexity of legal texts complicates the application of general augmentation techniques. Ensuring the quality of augmented data remains critical, with efforts focusing on preserving semantic and syntactic integrity to prevent models from learning incorrect patterns (Perçin et al., 2022). Additionally, the effectiveness of augmentation varies across models, necessitating tailored strategies, while scalability concerns arise with computationally intensive methods (Ghosh et al., 2023). The predominant focus on English limits applicability to multilingual legal contexts and in extremely low-resource scenarios, there is a risk of overfitting to augmented data (Ghosh et al., 2023). Future research should prioritise developing domain-specific augmentation techniques that maintain the nuances of legal text, explore flexible approaches, enhance scalability, and extend support diverse languages, thereby addressing these persistent challenges in legal NLP.
@@
In legal NLP, enhancing multilingual capabilities remains an underdeveloped area. While efforts, such as MultiLegalPile (Niklaus et al., 2024) have begun to address this, there remains a gap in research for many languages, including, but not limited to, Persian and Arabic. These limitations restrict the application of legal NLP across diverse legal systems worldwide, hampering broader adoption and accessibility. Multilingual capabilities introduce unique challenges for legal NLP models, primarily due to the distinct linguistic structures of each language, which often require extensive fine-tuning to ensure accuracy and relevancy in legal contexts. Furthermore, each legal system possesses its own set of terms and document standards, which can vary dramatically from one language to another. Therefore, expanding research into these and other underserved languages is essential for making NLP tools universally applicable.
ORC7: Ontologies and Knowledge Graphs
The use of ontologies in the legal domain is relatively sparse, yet it holds considerable potential to enhance the robustness of AI methodologies. Ontologies or knowledge graphs can also enable AI models to draw accurate inferences regarding the relationships between entities, thereby improving reasoning and decision making over complicated legal texts. However, utilising ontologies in legal NLP faces unique challenges. The complexity of legal language and the concept of “open texture,” where the meaning of legal terms can evolve over time, can complicate the creation of static ontological models (Mommers, 2010). Legal ontologies must be dynamic, reflecting changes in law and its interpretation over time. Additionally, the integration of real-world and legal concepts within ontologies presents further complexity, as it requires accommodating both legal terms and their relevant real-world contexts (Mommers, 2010).